40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 1 16:23:48 2018, Koji, Summary, Computers, Transition of the main NFS disk on chiara 
    Reply  Wed Aug 1 19:41:15 2018, gautam, Summary, Computers, RTCDS recovery, c1ioo changes CDS_overview.png
       Reply  Wed Aug 1 20:44:57 2018, gautam, Summary, Computers, c1omc model (re?)created CDS_OVERVIEW_withOMC.png
          Reply  Thu Aug 2 20:54:18 2018, gautam, Summary, Computers, c1omc model looks stable 
       Reply  Thu Aug 2 20:47:29 2018, gautam, Summary, Electronics, X Green "Mystery" solved 
          Reply  Thu Aug 2 23:09:25 2018, rana, Summary, Computers, X Green "Mystery" solved 
          Reply  Fri Aug 3 14:35:56 2018, gautam, Summary, Electronics, EX AUX electronics power restored 
    Reply  Mon Aug 6 09:42:10 2018, Koji, Summary, Computers, Transition of the main NFS disk on chiara 
Message ID: 14128     Entry time: Fri Aug 3 14:35:56 2018     In reply to: 14125
Author: gautam 
Type: Summary 
Category: Electronics 
Subject: EX AUX electronics power restored 

Steve and I restored the power to the EX AUX electronics rack. The power strip on the lowest shelf of the AUX rack now goes to another power strip laid out vertically along the NW corner of 1X9. The EX green locks to the arm just fine now.

ELOG V3.1.3-